OpenStudy (anonymous):

Philip has $2,000 and spends $23.75 on supplies. He divides the remaining amount equally among his 8 employees. How much does each employee receive?

OpenStudy (anonymous):

Initial Amount = 2000 Amount spent = 23.75 => Amount left = 2000-23.75 = 1976.25 No. of employees = 8 => Amount each employee gets = Total/No.of employees = 1976.25/8 = $247.03